The ratio of radii of gyration of a spherical shell and a solid sphere of the same mass and radius about a tangential axis is....pls explain the method

Grade:11

1 Answers

Khimraj
5 years ago
Radius of gyration of spherical shell
K1 = \sqrt{I/M}
I =  (2/3)MR2 + MR2 = (5/3)MR2
so K1 = \sqrt{5/3} R
Radius of gyration of solid sphere
K2 = \sqrt{I/M}
I =  (2/5)MR2 + MR2 = (7/5)MR2
so K2 = \sqrt{7/5} R
so K1/K2 = \sqrt{25/21}
Hope it clears............................
 
23𝑚𝑅2+𝑚𝐿2